Suppose you are given the following data for this company for the year:

                                                Price per unit                           $15

                                                Quantity sold 40,000 units

                                                Accounting depreciation         $20,000

                                                Forgone job at Firm A                        $80,000

                                                Economic depreciation            $45,000

                                                Utilities expense                     $60,000

                                                Wages paid                             $90,000

                                                Forgone job at Firm B            $90,000

                                                Money invested in business    $100,000

                                                Normal profit                          $14,000

                                                Materials expense                   $30,000

                                                Rental expense on factory       $140,000

                                                Current interest rate                5 percent

1. Calculate accounting profit for this firm for the year. Show your work.

2. Calculate economic profit for this firm for the year. Show your work.
Answer #1

(1)

Accounting profit ($) = Revenue - Explicit costs = (Price x Quantity) - Material expense - Wages paid - Utilities expense - Rental expense - Accounting depreciation - Normal profit

= (15 x 40,000) - 30,000 - 90,000 - 60,000 - 140,000 - 20,000 - 14,000

= 600,000 - 354,000

= 246,000

(2)

Economic profit ($) = Accounting profit - Economic depreciation - Interest income foregone - Maximum amount of salary foregone

= 246,000 - 45,000 - (100,000 x 5%) - Max(80000, 90000)

= 201,000 - 5,000 - 90,000

= 106,000